OpenStudy (anonymous):

The value, V(m), of a comic book m months after publication has a average rate of change of -0.04 between m = 36 and m = 60. Which statement must be true? A.) The value of the comic book decreased by a total of $0.04 between m = 36 and m = 60. B.) The value of the comic book increased by a total of $0.04 between m = 36 and m = 60. C.) The value of the comic book decreased by an average of $0.04 each month between m = 36 and m = 60. D.) The value of the comic book increased by an average of $0.04 each month between m = 36 and m = 60.

myininaya (myininaya):

So average rate of change from m=a to m=b can be found from the following formula: \[\frac{f(b)-f(a)}{b-a}\]

myininaya (myininaya):

you can replace f with V

myininaya (myininaya):

Didn't see the choices just a sec ago.

myininaya (myininaya):

But anyways the key word is average and you should also notice the sign that is in front of this average.

myininaya (myininaya):

IF \[\frac{V(b)-V(a)}{b-a}=\text{ negative number }\] then don't you think the average would be decreasing?

myininaya (myininaya):

and if we had \[\frac{V(b)-V(a)}{b-a}=\text{ positive number } \] then wouldn't the average be increasing?

myininaya (myininaya):

Think about the slope of lines. Because the average rate of change is using a secant line. IF the slope of the line is positive, then the line is increasing. If the slope of the line is negative, then the line is decreasing.